5) Nearly free electron model: We have seen that (6.43) gives an approximate solution of the Schrödinger for a weak potentialUck 0 Ck-g - E 2m Uck 0 Uck-g -E Ck 2m e (6.43) h2(k +g) 2me Uc 0. - E Ck+g
Answer #1

when introduced Energy of a particle in a periodic system we start from free particles ( w o ) and a small periodic potentialso ver va IGE since. V is real ie. von = V(o) we can also write - VerEVGene = V+ 2 Va Cos(678) [va oval 0 GO Gto Now for freeNow, K4 114&> ľ ERK 5 Vo katika kesi estos > G7O ktk where, hk zm 20 on. Keles, Ticiasclay: Va fi cick-wit mo da Vook, ktG GElk) : EC + E 1 € = tik? 2W 9th ---- G:29 In both case. E blows up. а The reason can not is apply b = Int are non degenerate147 = - a1k> + STK> alus + blk+G) Hotva)] 147 €147 La Elk> + av(x)16) + b Extalk+G) + b (x) 1k+G) = at lk) + BEIK+G) Takingplease thumb up